OpenStudy (anonymous):

Find the solution set to each equation. (2/x+1) - (3/x-1) = (5x+1/x^2-1)

OpenStudy (anonymous):

x=1 The detailed answer is given below. 2/(x+1)-3(x-1)=5x+1/(x^2-1) Or, 2(x-1) - 3(x+1)/(x^2-1)=5x+1/(x^2-1) Or, 2x-2-3x-3=5x+1 Or, -x-5=5x+1 Or, -6x=-6 Hence, x=1